21. The graph of ln A versus time for a reaction:

a. allows to determine if the reaction is exothermic

b. allows to calculate the value of the reaction rate constant

C. allows to calculate the activation energy of the reaction

d. all of the above
Answer #1

The relation between lnA and Actication energy Ea and rate constant k are related by Arrhenius equation as

lnk = lnA -(Ea/RT)

=> lnA = (Ea/RT) + lnk.....(1)

This follows the equation y = mx + c, so a plot of y vs x will give a straight line with slope m and intercept c.

Hence, above equation (1) while plotting lnA vs (1/T) will give a straight line with slope (Ea/R) and intercept lnk. From the slope we can calculate Ea, the activation energy of the reaction. Hence, the correct option is (C).